- Let A 1 5 3 1 -2 4 Find the element of W = Col(A) that is closest to 4 =-2 ܝ ܟܐ ܘܬ
Answer #1

The column space of A is the plane 1 W = span پا 5 1 4 -2 . The closest vector w\in W to y is the orthogonal projection of y onto W . That is,

у w

w = \text{proj}_{W} (y)
= y- \text{proj}_{W^\perp} ( y )
=y- \text{proj}_{n} (y)

where n is a normal vector to W . For example,

n = \begin{bmatrix} 1\\ 3 \\-2 \end{bmatrix}\times\begin{bmatrix} 5\\ 1 \\4 \end{bmatrix}
= \begin{vmatrix} \mathbf i & \mathbf j& \mathbf k\\ 1 & 3 & -2\\ 5 & 1 & 4 \end{vmatrix}
=\begin{bmatrix} 14 \\ -14 \\ -14\end{bmatrix}

since the constant 14 does not matter, we instead define

n =\begin{bmatrix} 1 \\ -1 \\ -1\end{bmatrix}.

Therefore,

w=y- \text{proj}_{n} (y)
= y - \frac{n\cdot y}{\|n\|^2} n
= \begin{bmatrix} 4 \\ -2 \\ -3 \end{bmatrix} - \frac{ \begin{bmatrix} 1 & -1 & -1 \end{bmatrix} \cdot \begin{bmatrix} 4 & -2 & -3 \end{bmatrix} }{1^2 + (-1^2)+ (-1)^2} \begin{bmatrix} 1 \\ -1 \\ -1 \end{bmatrix}
= \begin{bmatrix} 4 \\ -2 \\ -3 \end{bmatrix} - \frac{ 4 +2 + 3 }{3} \begin{bmatrix} 1 \\ -1 \\ -1 \end{bmatrix}
= \begin{bmatrix} 4 \\ -2 \\ -3 \end{bmatrix} - 3 \begin{bmatrix} 1 \\ -1 \\ -1 \end{bmatrix}
= \begin{bmatrix} 1 \\ 1 \\ 0 \end{bmatrix}
= 2\begin{bmatrix} 1 \\ 3 \\ -2 \end{bmatrix} +\begin{bmatrix} 5 \\ 1 \\ 4 \end{bmatrix} \in\text{col}(A) .